Phantomdex:

Two cylinders are given. A small cylinder has a radius of 2 meters, and a larger cylinder has a radius of 5 meters. The larger cylinder is a dilation of the smaller cylinder. Part A: Determine the scale factor. Show or explain all the necessary steps. (2 points) Part B: If the volume of the smaller cylinder is 24π m3, show how the scale factor can be used to determine the volume of the larger cylinder. (2 points)

Phantomdex:

1 attachment
joebama:

do you need a formula sheet?

Phantomdex:

perhaps

joebama:

1 attachment
Phantomdex:

@joebama wrote:
ty!

joebama:

@phantomdex wrote:
@joebama wrote:
ty!
anytime! perks of being home schooled

surjithayer:

scale factor=5/2=2.5 volume of cylinder\[v=\pi r^2 h\] volume dilation=(2.5)^3 volume of larger cylinder\[=(2.5)^3\times 24 \pi =375 \pi ~m^3\]

Phantomdex:

@surjithayer wrote:
scale factor=5/2=2.5 volume of cylinder\[v=\pi r^2 h\] volume dilation=(2.5)^3 volume of larger cylinder\[=(2.5)^3\times 24 \pi =375 \pi ~m^3\]
Ty